A daredevil on a motorcycle leaves the end of a ramp with a speed of 35.0 m/s .If his speed is 33.0 m/s when he reaches the peak of the path, what is the maximum height that he reaches? Ignore friction and air resistance.
1
Expert's answer
2020-03-09T11:15:18-0400

Let the angle at which the motorcycle leaves the ramp makes an angle "\theta" with the horizontal

the horizontal component= "35cos\theta"

vertical component = "35sin\theta"

At the maximum height the vertical component of the velocity becomes zero

"35cos\theta=33
\\cos\theta=\dfrac{33}{35}" "sin\theta=\dfrac{11.67}{35}"


vertical component ="35sin\theta =11.67m/s"

at maximum height vertical component of velocity is zero

So applying formula

"v^2=u^2+2gS"

"0=(11.67)^2+2(-10)S \\S=136/20=6.8 m"